# Effective Command-line Interface Fuzzing with Path-Aware Large Language Model Orchestration

### v1

@@ 1 Introduction
Fuzzing is a well-known and widely adopted approach to systematically discover software bugs and vulnerabilities, e.g., in Command-Line Interface (CLI) programs, with unexpected, malformed, or random inputs. Traditional fuzzing techniques, according to a survey of 102 recent fuzzing papers [lee2025zigzagfuzz], often operate with fixed option configurations, making them challenging to reach deeply-embedded vulnerable functions that are only triggerable with certain option strings. Even recent directed fuzzing [huang2022beacon, luo2023selectfuzz], which use distance metrics to guide mutations toward target functions, may face similar challenges because without a correct option combination, distance metric may stay at infinity regardless of other mutations.
Therefore, an important problem that CLI fuzzing approaches need to address is the generation of different combinations of option strings. Traditional CLI fuzzing approaches [wang2023carpetfuzz, song2020crfuzz, zhang2023fuzzing, lee2025zigzagfuzz] rely on CLI option parsers, which only capture syntactic knowledge but not semantic relationships between different options. ProphetFuzz [wang2024prophetfuzz], the only, state-of-the-art LLM-based CLI fuzzing approach, parses the man pages of CLI options using LLM to generate CLI option combinations that are likely to trigger vulnerabilities. However, sometimes, man pages do not contain the deep knowledge of different option combination either.
@@
We implemented a prototype of PILOT and plan to release it at https://github.com/momo-trip/PILOT-code. Our evaluation across 43 widely used programs shows that PILOT discovered 51 zero-day vulnerabilities. We responsibly disclosed all of them to their developers and so far, 41 have been confirmed by developers with 33 being fixed and three assigned CVE identifiers. We also compared PILOT with state-of-the-art CLI fuzzing approaches, and the results showed substantial improvements in code coverage, i.e., 16.0% over CarpetFuzz [wang2023carpetfuzz], 36.2% over ProphetFuzz [wang2024prophetfuzz], 21.8% over ZigZagFuzz [lee2025zigzagfuzz], and 57.9% over SelectFuzz [luo2023selectfuzz].
@@ 2.1 A Motivating Example
We describe a motivating example with a zero-day vulnerability found by PILOT in FFmpeg [ffmpeg], a multimedia framework for video and audio processing, which is widely used in popular applications like web browsers, media servers, and video conferencing systems. The vulnerability is a zero-day buffer overflow due to a negative size parameter, located in FFmpeg’s Real-time Transport Protocol (RTP) streaming functionality of mp4 files [ffmpeg_protocols]. It can be triggered either locally via command line or remotely through network streaming protocols, leading to arbitrary code execution through controlled heap/stack corruption. We have responsibly disclosed this vulnerability to the FFmpeg developers, who have acknowledged and fixed the vulnerability. CVE assignment is in progress.
@@ 2.2 Research Challenge and Overall Solution
To the best of our knowledge, state-of-the-art CLI fuzzing approaches [wang2023carpetfuzz, wang2024prophetfuzz, lee2025zigzagfuzz] have difficulties in triggering this vulnerability, because they cannot generate the seeds that satisfy or are close to the aforementioned two conditions.
@@ 3.4 Path-guided Context Prompting
Path-guided context prompting enhances LLM-based seed generation by providing explicit execution path information and enabling LLMs to understand the relationship between command-line arguments and target functions efficiently. This technique is complemented by an LLM-user interface that works together to overcome the inherent limitations of LLM context windows [window]. The following subsection describes the path-guided context prompting and the interface that enables autonomous code exploration.
@@ 3.6 Initial Seed Extraction from LLM-generated Test Suites
Extracting command-line options. Different mutators require seeds in specific formats. For example, CarpetFuzz [wang2023carpetfuzz] expects a single-line format: {command} [options] -i @@, where @@ is replaced by the fuzzer with mutated input files. Converting complex shell scripts into this format requires semantic understanding of the script structure. For example, consider the following shell script generated by PILOT:
@@ 4 Implementation and Experimental Setup
avconv [libav2021] 680,276 1,193 11,074 libav-git-c464278
bison [bison2021] 530,528 133 1,497 bison-3.7.6
cflow [cflow2021] 80,174 28 400 cflow-1.6
cjpeg [libjpegturbo2021] 112,689 39 312 libjpeg-turbo-2.1.0
dwarfdump [libdwarf2021] 183,051 109 1,503 libdwarf-20210528
ffmpeg [ffmpeg2021] 1,611,669 2,040 20,448 ffmpeg-N-103440
gm [graphicsmagick2021] 587,458 161 1,727 GraphicsMagick-1.3.36
gs [ghostscript2021] 3,527,477 1,497 21,673 ghostpdl-9.54.0
mpg123 [mpg123_2021] 113,461 70 733 mpg123-1.28.2
nasm [nasm2021] 158,205 79 832 nasm-2.15.05
objdump [binutils2021] 4,628,717 136 2,977 binutils-2.36.1
pspp [pspp2021] 732,585 435 4,765 pspp-1.4.1
readelf [binutils2021] 4,628,717 130 1,289 binutils-2.36.1
tiff2pdf [libtiff2021] 173,153 34 846 tiff-4.3.0
tiff2ps [libtiff2021] 173,153 38 867 tiff-4.3.0
vim [vim2021] 918,746 135 6,864 vim-8.2.3113
xmllint [libxml2_2021] 549,947 45 3,070 libxml2-2.9.12
xmlwf [expat2021] 67,906 10 417 expat-2.4.1
yara [yara2021] 62,743 64 603 yara-4.1.1
Implementation.   Our implementation contains 12,872 Lines of Python Code. The default LLM is based on Claude Sonnet 4.0 (claude-sonnet-4-20250514) [3_7_sonnet]. GPT-4.1 (gpt-4.1) [gpt] is also employed for the comparison. PILOT’s uses libclang [libclang] parser to generate call graphs, which is a C interface to the Clang compiler and provides access to Abstract Syntax Tree (AST) of the target C code. It also uses the NetworkX (nx) [NetworkX] library for graph analysis and centrality calculations.
Setup.   We conducted our experiments on Microsoft Azure virtual machines. Seed generation was performed on a Standard_L8as_v3 instance with 8 vCPUs, 64 GB RAM, and 512 GB storage running Ubuntu 22.04 LTS. Fuzzing campaigns were executed on a Standard_F32s_v2 instance with 32 vCPUs, 64 GB RAM, and 512 GB storage running Ubuntu 20.04 LTS. We set the maximum number of iterative refinements per function ($N_{trial}$) to 2 and the number of target functions ($N_{target}$) to 10. To reduce probabilistic behavior, the temperature [temperature] is set to 0, and max_tokens parameter [temperature], which indicates the maximum output tokens, is set to 4,096. All target programs were compiled with AddressSanitizer [serebryany2012addresssanitizer] enabled to detect memory safety violations including buffer overflows, use-after-free, and NULL pointer dereferences.
@@
(i) POWER Dataset (20 programs). We first included the 20 programs from the POWER [lee2022power] dataset, which has been widely used for command-line fuzzing evaluation in prior work including CarpetFuzz [wang2023carpetfuzz], ProphetFuzz [wang2024prophetfuzz], and ZigZagFuzz [lee2025zigzagfuzz]. These programs represent diverse applications ranging from parsers to multimedia tools and network utilities, with codebases between 62,743 and over 1,000,000 lines of code.
@@
CarpetFuzz [wang2023carpetfuzz] CarpetFuzz CarpetFuzz
ProphetFuzz [wang2024prophetfuzz] ProphetFuzz CarpetFuzz
@@
ZigZagFuzz [lee2025zigzagfuzz] (Default) dictionary ZigZagFuzz
@@
SelectFuzz [luo2023selectfuzz] No initial seeds SelectFuzz
@@
  • CarpetFuzz [wang2023carpetfuzz]: A state-of-the-art non-LLM approach that provides both initial seed generation and mutation capabilities. It leverages natural language processing to analyze command-line option inter-relationships from documentation for seed generation, and employs grammar-based mutations for fuzzing.
  • ProphetFuzz [wang2024prophetfuzz]: A state-of-the-art LLM-based approach for initial seed generation that leverages LLM’s knowledge on high-risk option combinations.
  • ZigZagFuzz [lee2025zigzagfuzz]: A state-of-the-art CLI fuzzing mutator that variates both option configurations and input files.
  • SelectFuzz [luo2023selectfuzz]: A state-of-the-art directed fuzzing approach that selectively guides exploration toward specific target locations by identifying and prioritizing relevant code paths.
@@
To enable CLI fuzzing across all approaches, we apply the following modifications. For ZigZagFuzz, while it can leverage built-in initial seed options of target programs, these options are program-specific and not universally applicable. We therefore prepare a default dictionary by extracting options from program help screens. For input files, we use the same default seed files from the fuzzing framework (AFL++’s basic seed corpus). For SelectFuzz, which by default lacks an interface for mutating option strings, we integrate AFL++’s argv fuzzing interface [aflplusplus_argv] into its fuzzer component.
@@ 5.1 RQ1: Zero-day Vulnerabilities
A critical challenge in reaching this function was that it only processes compiled YARA rule files (.yarac), which must be generated using the yarac compiler from source .yar files. Traditional seed generation approaches cannot orchestrate auxiliary tools like yarac to create the required input format. PILOT analyzed the call paths and identified this dependency, then generated a complete workflow: creating .yar rule files, compiling them with yarac to produce .yarac files, and invoking yara -C to load the compiled rules. PILOT selected this function as a target based on its high centrality in the call graph. Prior research [liu2020large] stated that high-centrality functions tend to contain more vulnerabilities, and this case validates that finding: the target function indeed contained a vulnerability.
@@ 5.3 RQ3: Coverage Improvement
The minimal set configuration represents the simplest possible invocation, containing only a single fixed option string, which is the typical approach used by many non-CLI-aware fuzzing frameworks [lee2025zigzagfuzz]. For input files, we use the AFL++ fuzzer’s default seed files for this minimal set configuration. Comparing against the minimal set (261 functions, 1,446 branches), we observe that CLI-aware approaches achieve substantially higher coverage. CarpetFuzz achieves 1.59× more functions, ProphetFuzz achieves 1.41×, and PILOT achieves 2.33×, demonstrating that systematic CLI option exploration is essential for reaching diverse program functionality.
@@ 6 Related Work
Command-line Fuzzing. AFLargv [Bohme2017Markov] extends AFL to support command-line argument fuzzing by processing command-line options using fixed-length data chunks. However, it applies traditional file-based fuzzers in a straightforward manner without modifying the core AFL fuzzing algorithm. TOFU [wang2020tofu] expands the fuzzing space to include command-line arguments and uses distance metrics to prioritize inputs that are closer to specified target basic blocks in the program. POWER [lee2022power] systematically explores command-line option configurations by constructing diverse option configurations and selecting maximally “distant” configurations based on function relevance. However, these approaches lack robust validity checking for generated option combinations, resulting in many invalid configurations that lead to early program termination.
@@
CrFuzz [song2020crfuzz] introduced clustering analysis to predict input validity, demonstrating improvements in path and edge coverage. ConfigFuzz [zhang2023fuzzing] implements a transformation technique that encodes program configurations as part of fuzzable input, allowing existing mutation operators to test program settings alongside normal inputs. CarpetFuzz [wang2023carpetfuzz] leverages natural language processing to automatically extract option constraints from documentation, significantly reducing the testing search space by filtering out invalid option combinations. ZigZagFuzz [lee2025zigzagfuzz] improves fuzzing coverage by separately mutating command-line options and file inputs in an interleaving manner. However, these studies do not prioritize option combinations based on their likelihood of improving coverage, leading to inefficient exploration of the vast option space.
@@
ProphetFuzz [wang2024prophetfuzz] uses LLMs to automatically predict high-risk command-line option combinations from program documentation. ProphetFuzz prioritizes combinations more likely to contain vulnerabilities and generates semantically coherent commands with appropriate input files. However, ProphetFuzz has two fundamental limitations. First, ProphetFuzz relies solely on user manuals and online documentation to infer CLI options, causing problems when documentation is incomplete, outdated, or insufficient for understanding option interactions. Second, ProphetFuzz generates input files through generic Python scripts rather than dedicated commands, which leads LLMs to produce ad-hoc files with limited validity and variety. In contrast, PILOT analyzes source code to obtain accurate option information and provides a sandboxed shell environment where the LLM can use real file-generation tools, enabling the generation of valid and diverse command-line options that reflect each program’s actual behavior and ultimately achieving substantially higher coverage.
@@
Directed Fuzzing. Directed fuzzing such as Beacon [huang2022beacon] and SelectFuzz [luo2023selectfuzz] guide fuzzing toward specific target locations. In particular, SelectFuzz employs a precondition-guided approach that selects seeds capable of satisfying branch conditions leading to target locations, improving the efficiency of reaching specific code regions. Other works such as Prospector [zhang2024prospector], Liu et al. [liu2020large], Cerebro [li2019cerebro], She et al.  [she2022effective], and Magneto [zhou2024magneto] have explored prioritization strategies and call chain decomposition for directed fuzzing. However, these techniques target specific code locations through traditional input fuzzing and do not address the unique challenges of CLI programs where command-line option configurations significantly determine reachable program paths and vulnerability exposure.
@@
LLM-based Test Generation. Numerous prior studies [yang2024enhancing, pizzorno2024coverup, dakhel2024effective, lemieux2023codamosa, xia2024fuzz4all] have proposed methods for generating test cases using LLMs. TELPA [yang2024enhancing] leverages program analysis to extract real usage scenarios and address hard-to-cover branches through feedback-based refinement. CoverUp [pizzorno2024coverup] employs coverage metrics to iteratively guide LLMs toward improved line and branch coverage, significantly outperforming prior methods. MuTAP [dakhel2024effective] utilizes mutation testing to identify weaknesses in generated tests and augment prompts with surviving mutants, enhancing bug detection capabilities. CITYWALK [zhang2025citywalk] specifically addresses C++’s complex features by incorporating project dependency analysis and language-specific knowledge to improve test correctness. These approaches demonstrate that combining structured program analysis with LLMs’ generative capabilities yields the effective automated testing tools, but they focus on unit testing and operate at the function level, making them unsuited for exercising CLI behavior.
@@
Fuzz4All [xia2024fuzz4all] leverages LLMs as input generation and mutation engines to produce diverse, realistic inputs for multiple programming languages. Oliinyk et al. [oliinyk2024fuzzing] presents research focused on embedded systems security, specifically targeting BusyBox, a ubiquitous software package in Linux-based embedded devices. However, these approaches rely on specific domain knowledge and cannot be universally applied to various programs.

### v2

# Iterative Large Language Model Prompting

@@ 1 Introduction
Fuzzing is a well-known and widely adopted approach to systematically discover software bugs and vulnerabilities, e.g., in Command-Line Interface (CLI) programs, with unexpected, malformed, or random inputs. Traditional fuzzing techniques, according to a survey of 102 recent fuzzing papers [22], often operate with fixed option configurations, making them challenging to reach deeply-embedded vulnerable functions that are only triggerable with certain option strings. Even recent directed fuzzing [20, 30], which use distance metrics to guide mutations toward target functions, may face similar challenges because without a correct option combination, distance metric may stay at infinity regardless of other mutations.
Therefore, an important problem that CLI fuzzing approaches need to address is the generation of different combinations of option strings. Traditional CLI fuzzing approaches [41, 39, 47, 22] rely on CLI option parsers, which only capture syntactic knowledge but not semantic relationships between different options. ProphetFuzz [42], the only, state-of-the-art LLM-based CLI fuzzing approach, parses the man pages of CLI options using LLM to generate CLI option combinations that are likely to trigger vulnerabilities. However, sometimes, man pages do not contain the deep knowledge of different option combination either.
@@
We implemented a prototype of PILOT and plan to release it at https://github.com/momo-trip/PILOT-code. Our evaluation across 43 widely used programs shows that PILOT discovered 51 zero-day vulnerabilities. We responsibly disclosed all of them to their developers and so far, 41 have been confirmed by developers with 33 being fixed and three assigned CVE identifiers. We also compared PILOT with state-of-the-art CLI fuzzing approaches, and the results showed substantial improvements in code coverage, i.e., 16.0% over CarpetFuzz [41], 36.2% over ProphetFuzz [42], 21.8% over ZigZagFuzz [22], and 57.9% over SelectFuzz [30].
@@ 2.1 A Motivating Example
We describe a motivating example with a zero-day vulnerability found by PILOT in FFmpeg [13], a multimedia framework for video and audio processing, which is widely used in popular applications like web browsers, media servers, and video conferencing systems. The vulnerability is a zero-day buffer overflow due to a negative size parameter, located in FFmpeg’s Real-time Transport Protocol (RTP) streaming functionality of mp4 files [12]. It can be triggered either locally via command line or remotely through network streaming protocols, leading to arbitrary code execution through controlled heap/stack corruption. We have responsibly disclosed this vulnerability to the FFmpeg developers, who have acknowledged and fixed the vulnerability. CVE assignment is in progress.
@@ 2.2 Research Challenge and Overall Solution
To the best of our knowledge, state-of-the-art CLI fuzzing approaches [41, 42, 22] have difficulties in triggering this vulnerability, because they cannot generate the seeds that satisfy or are close to the aforementioned two conditions.
@@ 3.4 Path-guided Context Prompting
Path-guided context prompting enhances LLM-based seed generation by providing explicit execution path information and enabling LLMs to understand the relationship between command-line arguments and target functions efficiently. This technique is complemented by an LLM-user interface that works together to overcome the inherent limitations of LLM context windows [3]. The following subsection describes the path-guided context prompting and the interface that enables autonomous code exploration.
@@ 3.6 Initial Seed Extraction from LLM-generated Test Suites
Extracting command-line options. Different mutators require seeds in specific formats. For example, CarpetFuzz [41] expects a single-line format: {command} [options] -i @@, where @@ is replaced by the fuzzer with mutated input files. Converting complex shell scripts into this format requires semantic understanding of the script structure. For example, consider the following shell script generated by PILOT:
@@ 4 Implementation and Experimental Setup
avconv [25] 680,276 1,193 11,074 libav-git-c464278
bison [16] 530,528 133 1,497 bison-3.7.6
cflow [17] 80,174 28 400 cflow-1.6
cjpeg [27] 112,689 39 312 libjpeg-turbo-2.1.0
dwarfdump [26] 183,051 109 1,503 libdwarf-20210528
ffmpeg [11] 1,611,669 2,040 20,448 ffmpeg-N-103440
gm [19] 587,458 161 1,727 GraphicsMagick-1.3.36
gs [6] 3,527,477 1,497 21,673 ghostpdl-9.54.0
mpg123 [32] 113,461 70 733 mpg123-1.28.2
nasm [33] 158,205 79 832 nasm-2.15.05
objdump [15] 4,628,717 136 2,977 binutils-2.36.1
pspp [18] 732,585 435 4,765 pspp-1.4.1
readelf [15] 4,628,717 130 1,289 binutils-2.36.1
tiff2pdf [28] 173,153 34 846 tiff-4.3.0
tiff2ps [28] 173,153 38 867 tiff-4.3.0
vim [31] 918,746 135 6,864 vim-8.2.3113
xmllint [14] 549,947 45 3,070 libxml2-2.9.12
xmlwf [10] 67,906 10 417 expat-2.4.1
yara [40] 62,743 64 603 yara-4.1.1
Implementation.   Our implementation contains 12,872 Lines of Python Code. The default LLM is based on Claude Sonnet 4.0 (claude-sonnet-4-20250514) [4]. GPT-4.1 (gpt-4.1) [2] is also employed for the comparison. PILOT’s uses libclang [8] parser to generate call graphs, which is a C interface to the Clang compiler and provides access to Abstract Syntax Tree (AST) of the target C code. It also uses the NetworkX (nx) [34] library for graph analysis and centrality calculations.
Setup.   We conducted our experiments on Microsoft Azure virtual machines. Seed generation was performed on a Standard_L8as_v3 instance with 8 vCPUs, 64 GB RAM, and 512 GB storage running Ubuntu 22.04 LTS. Fuzzing campaigns were executed on a Standard_F32s_v2 instance with 32 vCPUs, 64 GB RAM, and 512 GB storage running Ubuntu 20.04 LTS. We set the maximum number of iterative refinements per function ($N_{trial}$) to 2 and the number of target functions ($N_{target}$) to 10. To reduce probabilistic behavior, the temperature [5] is set to 0, and max_tokens parameter [5], which indicates the maximum output tokens, is set to 4,096. All target programs were compiled with AddressSanitizer [37] enabled to detect memory safety violations including buffer overflows, use-after-free, and NULL pointer dereferences.
@@
(i) POWER Dataset (20 programs). We first included the 20 programs from the POWER [21] dataset, which has been widely used for command-line fuzzing evaluation in prior work including CarpetFuzz [41], ProphetFuzz [42], and ZigZagFuzz [22]. These programs represent diverse applications ranging from parsers to multimedia tools and network utilities, with codebases between 62,743 and over 1,000,000 lines of code.
@@
CarpetFuzz [41] CarpetFuzz CarpetFuzz
ProphetFuzz [42] ProphetFuzz CarpetFuzz
@@
ZigZagFuzz [22] (Default) dictionary ZigZagFuzz
@@
SelectFuzz [30] No initial seeds SelectFuzz
@@
  • CarpetFuzz [41]: A state-of-the-art non-LLM approach that provides both initial seed generation and mutation capabilities. It leverages natural language processing to analyze command-line option inter-relationships from documentation for seed generation, and employs grammar-based mutations for fuzzing.
  • ProphetFuzz [42]: A state-of-the-art LLM-based approach for initial seed generation that leverages LLM’s knowledge on high-risk option combinations.
  • ZigZagFuzz [22]: A state-of-the-art CLI fuzzing mutator that variates both option configurations and input files.
  • SelectFuzz [30]: A state-of-the-art directed fuzzing approach that selectively guides exploration toward specific target locations by identifying and prioritizing relevant code paths.
@@
To enable CLI fuzzing across all approaches, we apply the following modifications. For ZigZagFuzz, while it can leverage built-in initial seed options of target programs, these options are program-specific and not universally applicable. We therefore prepare a default dictionary by extracting options from program help screens. For input files, we use the same default seed files from the fuzzing framework (AFL++’s basic seed corpus). For SelectFuzz, which by default lacks an interface for mutating option strings, we integrate AFL++’s argv fuzzing interface [1] into its fuzzer component.
@@ 5.1 RQ1: Zero-day Vulnerabilities
A critical challenge in reaching this function was that it only processes compiled YARA rule files (.yarac), which must be generated using the yarac compiler from source .yar files. Traditional seed generation approaches cannot orchestrate auxiliary tools like yarac to create the required input format. PILOT analyzed the call paths and identified this dependency, then generated a complete workflow: creating .yar rule files, compiling them with yarac to produce .yarac files, and invoking yara -C to load the compiled rules. PILOT selected this function as a target based on its high centrality in the call graph. Prior research [29] stated that high-centrality functions tend to contain more vulnerabilities, and this case validates that finding: the target function indeed contained a vulnerability.
@@ 5.3 RQ3: Coverage Improvement
The minimal set configuration represents the simplest possible invocation, containing only a single fixed option string, which is the typical approach used by many non-CLI-aware fuzzing frameworks [22]. For input files, we use the AFL++ fuzzer’s default seed files for this minimal set configuration. Comparing against the minimal set (261 functions, 1,446 branches), we observe that CLI-aware approaches achieve substantially higher coverage. CarpetFuzz achieves 1.59× more functions, ProphetFuzz achieves 1.41×, and PILOT achieves 2.33×, demonstrating that systematic CLI option exploration is essential for reaching diverse program functionality.
@@ 6 Related Work
Command-line Fuzzing. AFLargv [7] extends AFL to support command-line argument fuzzing by processing command-line options using fixed-length data chunks. However, it applies traditional file-based fuzzers in a straightforward manner without modifying the core AFL fuzzing algorithm. TOFU [43] expands the fuzzing space to include command-line arguments and uses distance metrics to prioritize inputs that are closer to specified target basic blocks in the program. POWER [21] systematically explores command-line option configurations by constructing diverse option configurations and selecting maximally “distant” configurations based on function relevance. However, these approaches lack robust validity checking for generated option combinations, resulting in many invalid configurations that lead to early program termination.
@@
CrFuzz [39] introduced clustering analysis to predict input validity, demonstrating improvements in path and edge coverage. ConfigFuzz [47] implements a transformation technique that encodes program configurations as part of fuzzable input, allowing existing mutation operators to test program settings alongside normal inputs. CarpetFuzz [41] leverages natural language processing to automatically extract option constraints from documentation, significantly reducing the testing search space by filtering out invalid option combinations. ZigZagFuzz [22] improves fuzzing coverage by separately mutating command-line options and file inputs in an interleaving manner. However, these studies do not prioritize option combinations based on their likelihood of improving coverage, leading to inefficient exploration of the vast option space.
@@
ProphetFuzz [42] uses LLMs to automatically predict high-risk command-line option combinations from program documentation. ProphetFuzz prioritizes combinations more likely to contain vulnerabilities and generates semantically coherent commands with appropriate input files. However, ProphetFuzz has two fundamental limitations. First, ProphetFuzz relies solely on user manuals and online documentation to infer CLI options, causing problems when documentation is incomplete, outdated, or insufficient for understanding option interactions. Second, ProphetFuzz generates input files through generic Python scripts rather than dedicated commands, which leads LLMs to produce ad-hoc files with limited validity and variety. In contrast, PILOT analyzes source code to obtain accurate option information and provides a sandboxed shell environment where the LLM can use real file-generation tools, enabling the generation of valid and diverse command-line options that reflect each program’s actual behavior and ultimately achieving substantially higher coverage.
@@
Directed Fuzzing. Directed fuzzing such as Beacon [20] and SelectFuzz [30] guide fuzzing toward specific target locations. In particular, SelectFuzz employs a precondition-guided approach that selects seeds capable of satisfying branch conditions leading to target locations, improving the efficiency of reaching specific code regions. Other works such as Prospector [48], Liu et al. [29], Cerebro [24], She et al.  [38], and Magneto [49] have explored prioritization strategies and call chain decomposition for directed fuzzing. However, these techniques target specific code locations through traditional input fuzzing and do not address the unique challenges of CLI programs where command-line option configurations significantly determine reachable program paths and vulnerability exposure.
@@
LLM-based Test Generation. Numerous prior studies [45, 36, 9, 23, 44] have proposed methods for generating test cases using LLMs. TELPA [45] leverages program analysis to extract real usage scenarios and address hard-to-cover branches through feedback-based refinement. CoverUp [36] employs coverage metrics to iteratively guide LLMs toward improved line and branch coverage, significantly outperforming prior methods. MuTAP [9] utilizes mutation testing to identify weaknesses in generated tests and augment prompts with surviving mutants, enhancing bug detection capabilities. CITYWALK [46] specifically addresses C++’s complex features by incorporating project dependency analysis and language-specific knowledge to improve test correctness. These approaches demonstrate that combining structured program analysis with LLMs’ generative capabilities yields the effective automated testing tools, but they focus on unit testing and operate at the function level, making them unsuited for exercising CLI behavior.
@@
Fuzz4All [44] leverages LLMs as input generation and mutation engines to produce diverse, realistic inputs for multiple programming languages. Oliinyk et al. [35] presents research focused on embedded systems security, specifically targeting BusyBox, a ubiquitous software package in Linux-based embedded devices. However, these approaches rely on specific domain knowledge and cannot be universally applied to various programs.